# Runtime Analysis of the SMS-EMOA for Many-Objective Optimization

### 16 Dec 2023 v1

@@ 1 Introduction
The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA). Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA. Zheng, Liu, and Doerr [ZLD22] conducted the first runtime analysis of the NSGA-II (see [ZD23a] for the journal version). This work quickly inspired many interesting follow-up results in bi-objective optimization [ZD22, BQ22, DQ23a, DQ23b, DQ23c, DOSS23a, DOSS23b, ${}^{+}$23]. In contrast to these positive results for two objectives, Zheng and Doerr [ZD23b] proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal. They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
@@
A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem [WD23]. Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
@@
With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by eume, Naujoks, and Emmerich [BNE07]. This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution. Many empirical works (see the almost 2,000 papers citing [BNE07]) confirmed the good performance of the SMS-EMOA for many-objective optimization. The only mathematical runtime analysis of the SMS-EMOA was conducted only very recently by Bian et al. [BZLQ23]. They proved that its runtime for the bi-objective OJZJ problem is $O(n^{k+1})$ and proposed a stochastic population update mechanism which was proven to achieve an expected runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$.
@@
Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives. We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem [DZ21], which is the problem analyzed in the only previous runtime analysis for the SMS-EMOA [BZLQ23]. We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge. We prove that the SMS-EMOA covers the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem), $m$ the number of objectives, and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
@@
We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$. Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in [WD23] suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
@@
We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update [BZLQ23] extends to the $m$-objective $m$OJZJ problem. Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
@@
On the positive side, we show that the advantage of heavy-tailed mutation is preserved. We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved. This is the same speed-up as observed for single-objective and bi-objective Jump problems [DLMN17, DZ21, ZD23c]. We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
@@
Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem [BZLQ23], we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ. We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ. These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
This paper is organized as follows. Section 2 briefly introduces some basic concepts for the many-objective optimization and the mathematical notations that will be used in this paper. The many-objective OJZJ benchmark is proposed in Section 3. Sections 4 to 6 analyze the runtime of the original SMS-EMOA, the SMS-EMOA with stochastic population update, and the one with heavy-tailed mutation on this many-objective benchmark. The runtime for two well-known bi-objective functions is conducted in Section 7. Section 8 concludes this work.
2 Preliminaries
@@
In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives. This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$. Here $m>2$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size. As in bi-objective optimization (that is, the case $m=2$), usually not all solutions are comparable. We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$, denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict. For a given many-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ w.r.t. $f$. The set of all Pareto optima is called Pareto set, and the set of all function values of the Pareto optima is called Pareto front. The goal for an MOEA is to find the full Pareto front as much as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front. As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front. We refer to [NW10, AD11, Jan13, ZYQ19, DN20] for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to [Bro11] for a discussion of runtime analyses of MOEAs.
@@
3 The $m$-Objective Jump Function
As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems [ZD23b]. In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed. Since the only runtime work of the SMS-EMOA so far [BZLQ23] analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well. Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem. With this, our results are comparable both the ones in [ZD23b] and [BZLQ23].
3.1 $m$OJZJ
We first recall the definition of the $m$OneMinMax problem. For the ease of presentation, we only consider even numbers $m$ of objectives here. In the $m$OneMinMax problem, the bit string (of length $n$) is divided into $m/2$ blocks of equal length $2n/m$. On each of these, a bi-objective OneMinMax problem is defined. We note that this general block construction goes back to the seminal paper of [LTZ04].
Definition 1 ([ZD23b]).
@@ 3.1 $m$OJZJ
We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark. It was first defined by Droste, Jansen, and Wegener [DJW02] and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima [JW02, ${}^{+}$18, Doe21, BBD21, HS22, RW22, Wit23, DEJK24].
@@
3.2 Characteristics
We now give more details on this $m$OJZJ function. Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block in the partition of $n$ bit positions. From Definition 2, we know that the bit values in the block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$. Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block. Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in [DZ21].
@@ 3.2 Characteristics
Note that for $k>1$, not all points are Pareto optimal and not all sets of mutually non-dominated points are the subsets of the Pareto front. However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front.
@@
Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominating (w.r.t. $m$OJZJ). Then $|P|\leq M$.
Proof.
Note that all objective values of any non-Pareto optimal solution are less than $2k$. Since each Pareto optimum has its all objective values of at least $2k$, we know that any non-Pareto optimal solution is dominated by all Pareto optima excluding the ones with some block all zeros or all ones (called extreme-block Pareto optima). Hence, $P$ excluding extreme-block Pareto optima, denoted as $P^{\prime}$, cannot include both Pareto optima and non-Pareto optimal solutions. If $P$ includes only Pareto optima, then we know that $|P|\leq M$.
We now consider that $P^{\prime}$ includes only non-Pareto optimal solutions, that is, for any $x\in P^{\prime}$, there exists a block $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$ such that $|x_{B_{i}}|_{1}\in[1..k-1]\cup[n^{\prime}-k+1..n^{\prime}-1]$. Note that for any block $B_{i}$, any two solutions $x,y$ are comparable w.r.t. $(f_{2i-1},f_{2i})$ if $(|x_{B_{i}}|_{1}\in[1..k-1],|y_{B_{i}}|_{1}\in[1..k-1])$ or $(|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1],|y_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$. Hence, with the bit values on other blocks fixed, there are at most two solutions in $P^{\prime}$ that are not weakly dominated. We also note that $(f_{2i-1}(z),f_{2i}(z))>(f_{2i-1}(x),f_{2i}(x))$ holds if $(z_{B_{i}}=1^{n^{\prime}},|x_{B_{i}}|_{1}\in[1..k-1])$ or $(z_{B_{i}}=0^{n^{\prime}},|x_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}-1])$. Hence, with the bit values on other blocks fixed, there are at most two solutions in $P$ that are not weakly dominated. Then $m/2$ blocks results in at most $2^{m/2}$ non-weakly-dominated solutions, that is, $|P|\leq 2^{m/2}<3^{m/2}\leq M$. ∎
@@
4 SMS-EMOA Can Optimize $m$OJZJ
As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of the $m$OneMinMax [ZD23b], which is a special case of $m$OJZJ. Motivated by the question whether there is a simple efficient variant of NSGA-II for many objectives, this section will analyze the runtime of the SMS-EMOA, which still uses the non-dominated sorting but replaces the crowding distance by hypervolume measure.
4.1 Algorithm Description
The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II. It maintains a fixed population size $\mu$. For each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$. From the $\mu+1$ individuals in the combined population $R_{t}=P_{t}\cup\{x^{\prime}\}$, one individual will be removed. It first sorts the combined population into several fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all mutually non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$. This is called non-dominated sorting. Different from the original NSGA-II that uses the crowding distance as the secondary criterion to remove the individual, SMS-EMOA removes the one with the smallest hypervolume contribution in the critical front, that is, $F_{i^{*}}$. The hypervolume of the set $S$ w.r.t. a reference point $r$ is defined as
$\operatorname{HV}_{r}(S)=\mathcal{L}\mathopen{}\mathclose{{}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}}\right),$
where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated by
@@ 4.1 Algorithm Description
Algorithm 1 shows the general procedure of the SMS-EMOA.
@@
5:   %Original Population Update
6:   Use fast-non-dominated-sort() [DPAM02] to divide $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
7:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
8:   Uniformly at random pick $z^{\prime}\in D$ and $P_{t+1}=R_{t}\setminus\{z^{\prime}\}$
9:end for
@@
4.2 Runtime of SMS-EMOA
We now establish the runtime of the SMS-EMOA to cover the full Pareto front of the $m$OJZJ. We first give the following lemma to show that for a large enough population size, the function value of any point with rank one in the combined population will survive to the next population. The key of the proof is that all $F_{1}$ points will survive if there is a point with a larger rank, and that at least two points have the same function value if all points are in $F_{1}$ and the population size is large enough.
@@ 4.2 Runtime of SMS-EMOA
Consider using SMS-EMOA with $\mu\geq M$ to optimize the $m$OJZJ problem. For any $x\in F_{1}$ for the current generation $t$, we have that there is $y$ with $f(y)=f(x)$ in $P_{t+1}$. Specifically, any Pareto front point will be maintained in all future populations once it is reached.
Proof.
Let $x\in F_{1}$. If $\cup_{i\geq 2}F_{i}\neq\emptyset$, then the removal will happen in $\cup_{i\geq 2}F_{i}$, and $x\in P^{t+1}$.
If all solutions are in $F_{1}$, from Lemma 3, we know that $|f(F_{1})|\leq M$. Hence, there exist $z_{1},z_{2}\in F_{1}$ such that $f(z_{1})=f(z_{2})$. Let $S_{x}=\{z\in R_{t}\mid f(x)=f(z)\}$. Then $|S_{x}|\geq 1$. If $|S_{x}|=1$, then $\Delta(x,F_{1})>0=\Delta(z_{1},F_{1})$, and thus $x\in P_{t+1}$. If $|S_{x}|>1$, since the survival selection only removes one solution, then there exists $y\in S_{x}$, such that $y\in P_{t+1}$.
Since any Pareto optimum is always in $F_{1}$, the latter part of this lemma holds. ∎
Recall that the inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before. Since not all points are inner Pareto optimum, the following lemma establishes the runtime to reach one inner Pareto optimum for the first time. The key of the proof is that at most $k$ specific bit flipping of a point that is not an inner Pareto optimum will result in an inner Pareto optimum.

### 25 Jan 2024 v2

@@ Introduction
The NSGA-II is the most widely-applied multiobjective evolutionary algorithm (MOEA). Non-dominated sorting and crowding distance are its two major features differentiating it from basic MOEAs such as the GSEMO or the $(\mu+1)$ SIBEA. Zheng, Liu, and Doerr (2022) conducted the first runtime analysis of the NSGA-II (see (Zheng and Doerr 2023a) for the journal version). This work quickly inspired many interesting follow-up results in bi-objective optimization (Zheng and Doerr 2022; Bian and Qian 2022; Doerr and Qu 2023a, b, c; Dang et al. 2023a, b; Cerf et al. 2023). In contrast to these positive results for two objectives, Zheng and Doerr (2023b) proved that for $m\geq 3$ objectives the NSGA-II needs at least exponential time (in expectation and with high probability) to cover the full Pareto front of the $m$-objective OneMinMax benchmark, a simple many-objective version of the basic OneMax problem where all search points are Pareto optimal. They claimed that the main reason for this low efficiency is the independent computation of the crowding distance in each objective.
@@
A very recent work showed that the NSGA-III, a successor algorithm of the NSGA-II aimed at better coping with many objectives, can efficiently solve the $3$-objective OneMinMax problem (Wietheger and Doerr 2023). Since apparently practitioners much prefer the NSGA-II (more than 17,000 citations on Google scholar only in the last five years) over the NSGA-III (5,035 citations since its publication in 2013), it remains an interesting question whether there are variants of the NSGA-II which better cope with many objectives.
@@
With the SMS-EMOA, an interesting variant of the NSGA-II was proposed by Beume, Naujoks, and Emmerich (2007). This algorithm is a steady-state variant of the NSGA-II (that is, in each iteration only a single offspring is generated and possibly integrated into the population) that further replaces the crowding distance as secondary selection criterion with the classic hypervolume contribution. Many empirical works (see the almost 2,000 papers citing (Beume, Naujoks, and Emmerich 2007)) confirmed the good performance of the SMS-EMOA for many-objective optimization. The first mathematical runtime analysis of the SMS-EMOA was conducted very recently by Bian et al. (2023), who proved that its expected runtime on the bi-objective OJZJ problem is $O(n^{k+1})$. They also proposed a stochastic population update mechanism and proved that it has the often superior runtime of $O(n^{k+1}\min\{1,n/2^{k/4}\})$. Zheng et al. (2024) proved that the SMS-EMOA has an expected runtime of $O(n^{4})$ on the bi-objective DLTB problem.
@@
Our Contributions: This paper conducts the first mathematical runtime analysis of the SMS-EMOA for more than two objectives. We first define the $m$OJZJ benchmark, an $m$-objective counterpart of the bi-objective OJZJ problem (Zheng and Doerr 2023d), which is the problem analyzed in the first runtime analysis for the SMS-EMOA (Bian et al. 2023). We note that the $m$OJZJ problem is the first multimodal many-objective benchmark proposed for a theoretical analysis, to the best of our knowledge. We prove that the SMS-EMOA covers the full Pareto front of this benchmark in an expected number of $O(M^{2}n^{k})$ iterations, where $n$ is the problems size, $k$ the gap size (a difficulty parameter of the problem), $m$ the number of objectives, and $M=(2n/m-2k+3)^{m/2}$ the size of the Pareto front.
@@
We recall that the original NSGA-II needs at least exponential time to optimize the $m$OneMinMax problem, which is a special case of $m$OJZJ with gap size $k=1$. Since the SMS-EMOA employs non-dominated sorting, but replaces the crowding distance in the original NSGA-II by the hypervolume contribution, our result in a similar fashion as the analysis of the NSGA-III in (Wietheger and Doerr 2023) suggests that the general approach of the NSGA-II is suitable for many-objective optimization and that it is only the crowding distance as tie-breaker which is not appropriate for more than two objectives.
@@
We then analyze whether the better performance of the SMS-EMOA on the bi-objective OJZJ problem achieved via a new stochastic population update (Bian et al. 2023) extends to the $m$-objective $m$OJZJ problem. Unfortunately, we shall observe that only a speed-up of order $1/\Theta(\min\{Mk^{1/2}/2^{k/2},1\})$ is obtained, which is $\Theta(1)$ when $m$ is large, e.g., $m>k$.
@@
On the positive side, we show that the advantage of heavy-tailed mutation is preserved. We analyze the SMS-EMOA with heavy-tailed mutation on the $m$OJZJ benchmark and prove that a speed-up of order $k^{0.5+k-\beta}$ is achieved. This is the same speed-up as observed for single-objective and bi-objective Jump problems (Doerr et al. 2017; Zheng and Doerr 2023d). We note that this is the first theoretical work to support the usefulness of heavy-tailed mutation in many-objective optimization.
@@
Finally, since so far the performance of the SMS-EMOA was only analyzed on the bi-objective OJZJ problem (Bian et al. 2023), we conduct mathematical runtime analyses of the SMS-EMOA also on the two most prominent bi-objective benchmarks OneMinMax and LOTZ. We prove that the SMS-EMOA finds the Pareto fronts of these benchmarks in an expected number of iterations of at most $2e(n+1)n(\ln n+1)$ for OneMinMax and at most $2en^{2}(n+1)$ for LOTZ. These are the same asymptotic runtimes (in terms of fitness evaluations) as known for the GSEMO and the NSGA-II.
Preliminaries
@@
In many-objective optimization, one tries to find good solutions for a problem containing several, but at least three objectives. This work considers pseudo-Boolean maximization, hence our problem is described by a function $f=(f_{1},\dots,f_{m}):\{0,1\}^{n}\rightarrow\mathbb{R}^{m}$. Here $m>2$ is the number of objectives and $n\in\mathbb{N}$ is called the problem size. As in bi-objective optimization (that is, the case $m=2$), usually not all solutions are comparable. We say that a solution $x\in\{0,1\}^{n}$ dominates a solution $y$, denoted as $x\succ y$, if $f_{i}(x)\geq f_{i}(y)$ for all $i\in\{1,\dots,m\}$, and at least one of these inequalities is strict. For a given many-objective function $f$, we say that $x\in\{0,1\}^{n}$ is a Pareto optimum if and only if there is no $y\in\{0,1\}^{n}$ that dominates $x$ w.r.t. $f$. The set of all Pareto optima is called Pareto set, and the set of all function values of the Pareto optima is called Pareto front. The goal for an MOEA is to find the full Pareto front as much as possible, that is, to compute a not too large set $S$ of solutions such that $f(S)$ equals or approximates well the Pareto front. As common in the mathematical runtime analysis of MOEAs, we call the runtime of an algorithm the number of function evaluations until its population $P$ covers the full Pareto front, that is, $f(P)$ contains the Pareto front. We refer to (Neumann and Witt 2010; Auger and Doerr 2011; Jansen 2013; Zhou, Yu, and Qian 2019; Doerr and Neumann 2020) for general introductions to the mathematical runtime analysis of evolutionary algorithms, and to (Brockhoff 2011) for a discussion of runtime analyses of MOEAs.
@@
The $m$-Objective Jump Function
As mentioned earlier, the NSGA-II was proven to have enormous difficulties in optimizing many-objective problems (Zheng and Doerr 2023b). In that paper, the $m$-objective counterpart $m$OneMinMax of the bi-objective OneMinMax benchmark was proposed and analyzed. Since the first runtime work of the SMS-EMOA so far (Bian et al. 2023) analyzed the bi-objective jump problem OJZJ, whose special case with gap size $k=1$ is (essentially) the OneMinMax problem, we shall propose and work with an $m$-objective version of OJZJ as well. Again, its special case $k=1$ will be (essentially) equal to the $m$OneMinMax problem. With this, our results are comparable both the ones in (Zheng and Doerr 2023b) and (Bian et al. 2023).
$m$OJZJ
We first recall the definition of the $m$OneMinMax problem. For the ease of presentation, we only consider even numbers $m$ of objectives here. In the $m$OneMinMax problem, the bit string (of length $n$) is divided into $m/2$ blocks of equal length $2n/m$. On each of these, a bi-objective OneMinMax problem is defined. We note that this general block construction goes back to the seminal paper of (Laumanns, Thiele, and Zitzler 2004).
Definition 1 (Zheng and Doerr (2023b)).
@@
We note that the function $\textsc{Jump}_{n^{\prime},k}$ used in the definition above is the famous Jump benchmark. It was first defined by Droste, Jansen, and Wegener (2002) and has quickly become the most employed multimodal benchmark in the theory of randomized search heuristics, leading to many fundamental results on how these algorithms cope with local optima (Jansen and Wegener 2002; Dang et al. 2018; Doerr 2021; Benbaki, Benomar, and Doerr 2021; Hevia Fajardo and Sudholt 2022; Rajabi and Witt 2022; Witt 2023; Doerr et al. 2024).
@@
Characteristics
We now give more details on this $m$OJZJ function. Let $B_{i}:=[(i-1)n^{\prime}+1..in^{\prime}]$, $i\in[1..m/2]$, be the $i$-th block in the partition of $n$ bit positions. From Definition 2, we know that the bit values in the block $B_{i}$ only influence the objectives $f_{2i-1}$ and $f_{2i}$. Figure 1 plots the objective values of $f_{2i-1}$ and $f_{2i}$ relative to the number of ones in this block. Obviously, $m$OJZJ is multimodal with respect to the definition of multimodality of multiobjective problems in (Zheng and Doerr 2023d).
@@
Note that for $k>1$, not all points are Pareto optimal and not all sets of mutually non-dominated points are the subsets of the Pareto front. However, we have the following result showing that the maximum number of mutually non-dominated points is at most the size of the Pareto front. Due to the limited space, all mathematical proofs could only be sketched or had to be omitted completely. They can be found in the preprint (Zheng and Doerr 2023c).
@@
Let $k\leq n^{\prime}/2$ and $P$ be a set of solutions such that any two are not weakly dominated (w.r.t. $m$OJZJ). Then $|P|\leq M$.
@@
The SMS-EMOA Can Optimize $m$OJZJ
As discussed before, the original NSGA-II with non-dominated sorting and crowding distance needs at least exponential runtime to cover the full Pareto front of $m$OneMinMax, $m\geq 3$, (Zheng and Doerr 2023b), which is a special case of $m$OJZJ. In this section, we analyze the runtime of the SMS-EMOA on this problem and show that it does not encounter such problems.
Algorithm Description
The SMS-EMOA is a steady-state variant (that is, the offspring size is smaller than the parent size) of the NSGA-II. Like the NSGA-II, the SMS-EMOA works with a population of fixed size $\mu$. However, in each iteration, only one offspring $x^{\prime}$ is generated from the parent population $P_{t}$. From the $\mu+1$ individuals in the combined parent and offspring population $R_{t}=P_{t}\cup\{x^{\prime}\}$, a single individual is removed. To this aim, the SMS-EMOA like the NSGA-II uses non-dominated sorting, that is, it partitions $R_{t}$ into fronts $F_{1},\dots,F_{i^{*}}$, where $F_{i}$ contains all non-dominated individuals in $R_{t}\setminus(\bigcup_{j=1}^{i-1}F_{j})$. Different from the original NSGA-II that now uses the crowding distance as the secondary criterion for the removal of individuals, the SMS-EMOA removes the individual with the smallest hypervolume contribution in the critical front (which here always is the last front $F_{i^{*}}$).
The hypervolume of a set $S$ of individuals w.r.t. a reference point $r$ in the objective space is defined as
$\operatorname{HV}_{r}(S)=\mathcal{L}\left(\,\bigcup_{u\in S}\{h\in\mathbb{R}^{m}\mid r\leq h\leq f(u)\}\right),$
where $\mathcal{L}$ is the Lebesgue measure. The hypervolume contribution of an individual $x\in F_{i^{*}}$ is calculated via
@@
Algorithm 1 gives the pseudocode of the SMS-EMOA.
@@
5:   Use fast-non-dominated-sort() (Deb et al. 2002) to partition $R_{t}=P_{t}\cup\{x^{\prime}\}$ into $F_{1},\dots,F_{i^{*}}$
6:   Calculate $\Delta_{r}(z,F_{i^{*}})$ for all $z\in F_{i^{*}}$ and find $D=\arg\min_{z\in F_{i^{*}}}{\Delta_{r}(z,F_{i^{*}})}$
7:   Uniformly at random pick $z^{\prime}\in D$ and set $P_{t+1}=R_{t}\setminus\{z^{\prime}\}$
8:end for
@@
Runtime of SMS-EMOA
We now analyze the runtime of the SMS-EMOA, that is, the time until its population covers the full Pareto front of $m$OJZJ. We start by proving the following result ensuring that certain individuals survive in the population or are replaced by at least as good individuals. We formulate this result in manner more general than what we need since we expect this general version to be useful in other analyses of the SMS-EMOA.
The key to the proof is that (i) if there is a point with a rank larger than one, then all solutions with rank one (that is, in $F_{1}$) will survive, and (ii) if $F_{1}$ contains two or more solutions with same function value, then only such a solution can be removed from $F_{1}$. If the population size is large enough, by the pigeon-hole principle, one of these two cases comes true.
@@
Consider any $m$-objective optimization problem. Let $M\in\mathbb{N}$ be such that any set $S$ of pairwise non-dominated solutions having different objective values satisfies $|S|\leq M$. Consider solving this problem via the SMS-EMOA with population size $\mu\geq M$ and using a reference point $r$ such that $\operatorname{HV}_{r}(\{x\})>0$ for any individual $x\in\{0,1\}^{n}$.
Then the following is true. If at some time $t$ the combined parent and offspring population $R_{t}$ of the SMS-EMOA contains some solution $x$ (and thus in particular if $x\in P_{t}$), then at any later time $s>t$ the parent population $P_{s}$ contains a solution $y$ such that $y\succeq x$.
In particular, if $R_{t}$ contains a Pareto optimum $x$, then all future generations contain a solution $y$ with $f(y)=f(x)$.
Recall that an inner Pareto optimum is a Pareto optimum with $x_{B_{i}}\notin\{0^{n^{\prime}},1^{n^{\prime}}\}$ for all blocks $B_{i}=[(i-1)n^{\prime}+1..in^{\prime}],i\in[1..m/2]$, as defined before. While it is very likely that at least one initial solution is an inner Pareto optimum, this does not happen with probability one, and hence in the following lemma we estimate the time until the population contains an inner Pareto optimum. Since this time usually is much smaller than the time to generate all remaining Pareto optima, we do not care that this is estimate could easily be improved.
The key to the proof is to note that changing a block $x_{B_{i}}$ to a bit-string with between $k$ and $n^{\prime}-k$ ones is relatively easy. By Lemma 4, an individual with this $f_{i}$ value will remain in the population. Hence a total of at most $m$ such block changes (applied to the right individual) suffice to obtain an inner Pareto optimum.